# Computing Lower and Upper Hitting Probabilities for Imprecise Markov Chains

### v1

@@ 2.5 Imprecise Markov Chains
Instead of a single stochastic process $\mathbb{P}$, we now consider a set of stochastic processes $\mathcal{P}$. If these processes are Markov chains, it is said to be an imprecise Markov chain (IMC). An interpretation may be that we do not know precisely which $\mathbb{P}\in\mathcal{P}$ is the true process so that the set $\mathcal{P}$ serves as an “error margin”, and, we hope, contains the true process $\mathbb{P}$ being modelled as well. Inferences with respect to the imprecise Markov chain $\mathcal{P}$ are defined using lower and upper expectations, which are defined as
@@
These inferences are conservative, tight lower and upper bounds on inferences with respect to all stochastic processes in the set $\mathcal{P}$. Lower and upper probabilities, $\underline{\mathbb{P}}_{\mathcal{P}}$ and $\overline{\mathbb{P}}_{\mathcal{P}}$, are a special case of lower and upper expectations:
@@
The set $\mathcal{P}$ representing the IMC will be parametrised through a nonempty set of transition matrices $\mathcal{T}$. Here we will focus only on (time-)homogeneous Markov chains $\mathbb{P}_{T}\in\mathcal{P}$ that have their transition matrix $T\in\mathcal{T}$. This choice is justified by the fact that Krak et al. [krak2019hitting] showed that lower (and upper) hitting probabilities remain unchanged even when the set $\mathbb{P}$ is enlarged to include (time-)inhomogeneous Markov processes compatible with $\mathcal{T}$.
@@
$\underline{p}_{\mathcal{T}}\coloneqq\inf_{T\in\mathcal{T}}p^{T}\quad\text{and}\quad\overline{p}^{\mathcal{T}}\coloneqq\sup_{T\in\mathcal{T}}p^{T}.$
Hence, we have that
@@
For a (precise) homogeneous Markov chain, the corresponding transition matrix $T$ provides valuable computational tools. We introduce analogous objects for imprecise Markov chains: the lower and upper transition operators.
Definition 2.15 (Lower and Upper Transition Operators [DeCooman2009]).
An operator $\underline{T}$ (resp. $\overline{T}$) is a lower (resp. upper) transition operator if, for all $f,g\in\mathbb{R}^{\mathcal{X}}$, all $\lambda>0$, and all $x\in\mathcal{X}$, it satisfies
  1. 1.
  2. 2.
    $\underline{T}f+\underline{T}g\leq\underline{T}(f+g)$ and $\overline{T}f+\overline{T}g\geq\overline{T}(f+g),$
  3. 3.
    $\underline{T}(\lambda f)=\lambda\underline{T}f$ and $\overline{T}(\lambda f)=\lambda\overline{T}f.$
We note that lower and upper transition operators are conjugate, meaning that any lower transition operator $\underline{T}$ induces a corresponding upper transition operator
@@
and vice-versa. Furthermore, any transition matrix is a lower transition operator with itself as the conjugate upper transition operator. Therefore, all properties derived for lower and upper transition operators also hold for transition matrices.
Beyond the properties in Definition 2.15, transition operators satisfy some additional useful properties.
Lemma 2.16 ([krak2019hitting]).
Given any lower transition operator $\underline{T}$ and its conjugate upper transition operator $\overline{T}$, for all $f,g\in\mathbb{R}^{\mathcal{X}}$ and all $n\in\mathbb{N}$ it holds that $\underline{T}^{n}$ is a lower transition operator, $\overline{T}^{n}$ its conjugate, and
@@
Fix some nonempty set of transition matrices $\mathcal{T}$. We will define the lower and upper transition operators corresponding to $\mathcal{T}$ as the lower and upper envelopes of $\mathcal{T}$, respectively.
Definition 2.17 ([krak2019hitting]).
Given a nonempty set of transition matrices $\mathcal{T}$, the corresponding lower and upper transition operators $\underline{T}$ and $\overline{T}$ are defined as
for all $f\in\mathbb{R}^{\mathcal{X}}$ and all $x\in\mathcal{X}$.
Throughout the paper, we assume that $\mathcal{T}$ is nonempty, compact, convex, and with separately specified rows (SSR) [HermansSkulj2014, krak2019hitting]. Equivalently, $\mathcal{T}$ can be represented as $N$ compact and convex sets of probability distributions $\{\mathcal{T}_{x}\}_{x\in\mathcal{X}}$, one for each state $x\in\mathcal{X}$, so $\mathcal{T}$ is the Cartesian product of these sets. Each of these sets is often called credal set.
Proposition 2.18 ([krak2019hitting]).
Let $\mathcal{T}$ be a nonempty, compact, convex set of transition matrices with SSR on the space $\mathcal{X}$. Then, for all $f\in\mathbb{R}^{\mathcal{X}}$, there exists an extreme point $T\in\mathcal{T}$ such that $Tf=\underline{T}f$, and there exists an extreme point $T^{\prime}\in\mathcal{T}$ such that $T^{\prime}f=\overline{T}f$.
@@ 3 Characterization of Lower and Upper Hitting Probabilities
In this section, we characterise lower and upper hitting probabilities as the unique solutions of certain systems of equations. We first formalise the notion of reachability in the imprecise framework and use it to identify the states that can reach the target set with positive lower (resp. upper) probability.
@@ 3.1 Lower and Upper Reachability
Before characterising the lower and upper hitting probabilities, we identify the subset of states for which these probabilities are trivial, i.e. equal to zero or one. Recall that in the precise setting, for a homogeneous Markov chain with transition matrix $T$, we focused on the set of states $A^{c}\setminus\mathcal{C}_{T}$. This was done because Proposition 2.5 directly yields the hitting probabilities for those states, with the hitting probability equal to one for states in $A$ and equal to zero for states in $\mathcal{C}_{T}$. Clearly, the result for the states in $A$ carries over to the imprecise setting, as $p^{T}(x)=1$ for all $T\in\mathcal{T}$ and all $x\in A$. Hence also $\underline{p}^{\mathcal{T}}(x)=\overline{p}^{\mathcal{T}}(x)=1$ for all $x\in A$. To characterise the states with zero lower and upper hitting probability, we need suitable definitions of reachability for the imprecise setting.
Definition 3.1 (Upper reachability [DeCooman2009, hermans2012characterisation, debock2017imprecise]).
Given a set of transition matrices $\mathcal{T}$ we say that a set of states $\mathcal{D}\subseteq\mathcal{X}$ is upper reachable from $x\in\mathcal{X}$ if there exists $n\in\mathbb{N}_{0}$ such that $\overline{T}^{n}\mathbb{1}_{\mathcal{D}}](x)>0$, i.e. there exists $n\in\mathbb{N}_{0}$ and $T\in\mathcal{T}$ such that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$. We denote by $x\rightharpoonup\mathcal{D}$ if $\mathcal{D}$ is upper reachable from $x$.
The fact that $x\rightharpoonup\mathcal{D}$ if and only if there exists $T\in\mathcal{T}$ and $n\in\mathbb{N}_{0}$ such that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$, follows from the fact that $\overline{T}^{n}\mathbb{1}_{\mathcal{D}}](x)>0$ implies the existence of a sequence of matrices $T_{1},\dots,T_{n}\in\mathcal{T}$ such that
$T_{1}\cdots T_{n}\mathbb{1}_{\mathcal{D}}](x)>0.$
From this sequence, and using the fact that the set $\mathcal{T}$ is convex and has SSR, we construct the desired matrix $T$.
@@
Given a set of transition matrices $\mathcal{T}$ we say that a set of states $\mathcal{D}\subseteq\mathcal{X}$ is lower reachable from $x\in\mathcal{X}$ if for all $T\in\mathcal{T}$ there exists $n\in\mathbb{N}_{0}$ such that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$. We denote by $x\rightharpoondown\mathcal{D}$ if $\mathcal{D}$ is lower reachable from $x$.
The definition of upper reachability is widely accepted meanwhile the same does not hold for the lower reachability one. In fact, in the literature [debock2025convergent, hermans2012characterisation, sangalli2025meeting, debock2017imprecise, erreygers2023], at least three other definitions are used.
@@
and where $k^{*}$ is the first index for which $\mathcal{D}_{k^{*}}=\mathcal{D}_{k^{*}+1}$. [debock2017imprecise, erreygers2023]
@@
$x$ lower reaches $\mathcal{D}$ if there exists $n\in\mathbb{N}_{0}$ such that for all $T\in\mathcal{T}$ it holds that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$. [sangalli2025meeting]
@@
$x$ lower reaches $\mathcal{D}$ if $\underline{T}^{n}\mathbb{1}_{\mathcal{D}}](x)>0$ for some $n\in\mathbb{N}_{0}$. [debock2025convergent, sangalli2025meeting]
Let us note that one reason for this abundance of possible definitions of lower reachability is that the quantifiers do not commute, i.e. in general
@@
$\operatorname{(LR0)}\Rightarrow\operatorname{(LR1)}:$ Suppose ex-absurdo that $x\notin\mathcal{D}_{k^{*}}$. This implies that there exists $\tilde{T}\in\mathcal{T}$ such that $\tilde{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](x)=0$.
Let $x=x_{0},\dots,x_{n}=y\in\mathcal{D}$ be a simple path from Proposition 2.2 that connects $x$ to $y$ using $\tilde{T}$. This path exists because $x\rightharpoondown\mathcal{D}$ under $\operatorname{(LR0)}$. Since $\tilde{T}(x,z)=0$ for all $z\in\mathcal{D}_{k^{*}}$ and $\tilde{T}(x,x_{1})>0$, it follows that $x_{1}\notin\mathcal{D}_{k^{*}}$. Note that the condition $\tilde{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](x)=0$ involves only the $x$th row of the matrix $\tilde{T}$. Therefore, since $\mathcal{T}$ has SSR and $x_{1}\notin\mathcal{D}_{k^{*}}$, there exists a modification of the $x_{1}$th row of $\tilde{T}$ such that $\tilde{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](x_{1})=0$. We then proceed along the path: if $\tilde{T}(x_{1},x_{2})>0$ we find that $x_{2}\notin\mathcal{D}_{k^{*}}$, and so on. Thus, we either find $x_{j}\notin\mathcal{D}_{k^{*}}$, for all $0\leq j\leq n$, which is in contradiction with $x_{n}\in\mathcal{D}\subseteq\mathcal{D}_{k^{*}}$, or we find an index $j^{*}$ such that by imposing $\tilde{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](x_{j^{*}})=0$ we get $\tilde{T}(x_{j^{*}},x_{j^{*}+1})=0$. This means that $x_{j}\notin\mathcal{D}_{k^{*}}$, for all $0\leq j\leq j^{*}$, while $x_{j^{*}+1}$ is in $\mathcal{D}_{k^{*}}$ and, for that reason, $\tilde{T}(x_{j^{*}},x_{j^{*}+1})$ has been set to zero.
If $x_{j^{*}}\not\to^{\tilde{T}}\mathcal{D}$ with this modified $\tilde{T}$, we restart from $x$. By our hypothesis, definition $\operatorname{(LR0)}$, and Lemma 2.14, there exists $m\in\mathbb{N}$, a matrix $\hat{T}\in\mathcal{T}$, with $\hat{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](x_{j})=0$ for all $1\leq j\leq j^{*}$, and a path $x=y_{0},\dots,y_{m}\in\mathcal{D}$ such that $0<p^{\hat{T}}(x)\leq\dots\leq p^{\hat{T}}(y_{m})=1$. This path cannot contain $x_{j^{*}}$ since its hitting probability is zero. As before, we impose $\hat{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](y_{i})=0$ along the path until we either reach $\mathcal{D}$ and get a contradiction or we encounter the first state $y_{i^{*}}$ for which $\hat{T}(y_{i^{*}},y_{i^{*}+1})=0$. If $y_{i^{*}}$ cannot reach $\mathcal{D}$ we eliminate $y_{i^{*}}$ and restart from $x$: the finiteness of $\mathcal{X}$ guarantees that we may repeat this argument only a finite number of times.
If $y_{i^{*}}\to^{\hat{T}}\mathcal{D}$, let $y_{i^{*}}=w_{0},\dots,w_{l}\in\mathcal{D}$ be a simple path that connects $y_{i^{*}}$ to $\mathcal{D}$ using this modified version of $\hat{T}$. Again, we follow this path until we get a contradiction or we encounter the first index $h^{*}$ for which by imposing $\tilde{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](w_{h^{*}})=0$ we get $\tilde{T}(w_{h^{*}},w_{h^{*}+1})=0$. We argue that, since $w_{h^{*}+1}\in\mathcal{D}_{k^{*}}$, $w_{h^{*}}$ is not in $\{x_{0},\dots,x_{j^{*}}\}$ or $\{y_{0},\dots,y_{i^{*}}\}$. Indeed, for these states, we have already set $\tilde{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](\cdot)=0$, so the path $y_{i^{*}}=w_{0},\dots,w_{l}$ that included one of these states could not have had positive transition probabilities. Therefore, each time we build a path we discover a new state not in $\mathcal{D}_{k^{*}}$: using the finiteness of $\mathcal{X}$ we get a contradiction.
$\operatorname{(LR1)}\Rightarrow\operatorname{(LR0)}:$ If $x\in\mathcal{D}_{1}\setminus\mathcal{D}$ then, by the definition of the sequence $(\mathcal{D}_{k})$, $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$ for all $T\in\mathcal{T}$, with $n=1$.
We will now proceed by strong induction on the following proposition:
$(P_{k}):$
@@
If $x\in\mathcal{D}_{k}\setminus\mathcal{D}_{k-1}$, then for all $T\in\mathcal{T}$ there exists $n\leq k$ such that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$.
Suppose that for all $w\in\mathcal{D}_{j}\setminus\mathcal{D}_{j-1}$ and all $1\leq j<k$ we have that $(P_{j})$ is true, i.e. there exists a path of length $m\leq j$ connecting $w$ to $\mathcal{D}$. This implies that for all $w\in\mathcal{D}_{k-1}$ and for all $T\in\mathcal{T}$ there exists a path of length $m\leq k-1$ connecting $w$ to $\mathcal{D}$, i.e.
${T}^{m}\mathbb{1}_{\mathcal{D}}](w)=\sum_{\begin{subarray}{c}i=2,\dots,m\\ x_{i}\in\mathcal{X}\\ x_{m+1}\in\mathcal{D}\end{subarray}}{T}(w,x_{2})\cdots{T}(x_{m},x_{m+1})>0.$
For all $x\in\mathcal{D}_{k}\setminus\mathcal{D}_{k-1}$ and all $T\in\mathcal{T}$ it holds that
$T\mathbb{1}_{\mathcal{D}_{k-1}}](x)=\sum_{w\in\mathcal{D}_{k-1}}{T}(x,w)>0.$
It follows that for all $T\in\mathcal{T}$ there exists $m\leq k-1$ such that
$[{T}^{m+1}\mathbb{1}_{\mathcal{D}}](x)$ $=\sum_{\begin{subarray}{c}i=1,\dots,m\\ x_{i}\in\mathcal{X}\\ x_{m+1}\in\mathcal{D}\end{subarray}}{T}(x,x_{1})\cdots{T}(x_{m},x_{m+1})$
$\geq\sum_{w\in\mathcal{D}_{k-1}}{T}(x,w)\sum_{\begin{subarray}{c}i=2,\dots,m\\ x_{i}\in\mathcal{X}\\ x_{m+1}\in\mathcal{D}\end{subarray}}{T}(w,x_{2})\cdots{T}(x_{m},x_{m+1})$
$=\sum_{w\in\mathcal{D}_{k-1}}{T}(x,w)[{T}^{m}\mathbb{1}_{D}](w)>0,$
@@
In the rest of the paper we will use (LR0) (Definition 3.2) (or the equivalent (LR1)) to define lower reachability. (LR0) is weaker than (LR2) and (LR3), namely
$\text{(LR3)}\Rightarrow\text{(LR2)}\Rightarrow\text{(LR1)}\Leftrightarrow\text{(LR0)}.$ (11)

### v2

@@ 2.5 Imprecise Markov Chains
Instead of a single stochastic process $\mathbb{P}$, we now consider a set of stochastic processes $\mathcal{P}$. If these processes are Markov chains, it is said to be an imprecise Markov chain (IMC) [5, 14, 12]. An interpretation may be that we do not know precisely which $\mathbb{P}\in\mathcal{P}$ is the true process so that the set $\mathcal{P}$ serves as an “error margin” and, we hope, contains the true process $\mathbb{P}$ being modelled as well. Inferences with respect to the imprecise Markov chain $\mathcal{P}$ are defined using lower and upper expectations, which are defined as
@@
These inferences are conservative, tight lower and upper bounds on inferences with respect to all stochastic processes in the set $\mathcal{P}$. Lower and upper probabilities, $\underline{\mathbb{P}}_{\mathcal{P}}$ and $\overline{\mathbb{P}}_{\mathcal{P}}$ respectively, are a special case of lower and upper expectations:
@@
The set $\mathcal{P}$ representing the IMC will be parametrised through a nonempty set of transition matrices $\mathcal{T}$. In this paper, we focus on (time-)homogeneous Markov chains $\mathbb{P}_{T}\in\mathcal{P}$ that have their transition matrix $T\in\mathcal{T}$. This choice is justified by Krak et al. [19, Theorem 18] showing that lower (and upper) hitting probabilities remain unchanged even when the set $\mathcal{P}$ is enlarged to include (time-)inhomogeneous Markov processes compatible with $\mathcal{T}$.
@@
$\underline{p}^{\mathcal{T}}\coloneqq\inf_{T\in\mathcal{T}}p^{T}\quad\text{and}\quad\overline{p}^{\mathcal{T}}\coloneqq\sup_{T\in\mathcal{T}}p^{T},$
where the infimum and supremum are taken pointwise. Thus, we have
@@
Throughout the paper, we assume that the set $\mathcal{T}$ is nonempty, compact, convex, and with separately specified rows (SSR) [14, 19]. This last property means that $\mathcal{T}$ can be represented as $N=|\mathcal{X}|$ compact and convex sets of probability distributions $\{\mathcal{T}_{x}\}_{x\in\mathcal{X}}$, one for each state $x\in\mathcal{X}$, so that $\mathcal{T}$ is the Cartesian product of these sets. Each of these sets is often called credal set. For such a set of transition matrices $\mathcal{T}$, we introduce valuable computational tools in its associated lower and upper transition operators.
Definition 2.15 ([19]).
Given a set of transition matrices $\mathcal{T}$ on $\mathcal{X}$, the corresponding lower and upper transition operators $\underline{T},\overline{T}:\mathbb{R}^{\mathcal{X}}\to\mathbb{R}^{\mathcal{X}}$ are defined as
for all $f\in\mathbb{R}^{\mathcal{X}}$ and all $x\in\mathcal{X}$.
From this definition, it follows that
$\underline{T}f](x)=\underline{\mathbb{E}}_{\mathcal{P}_{\mathcal{T}}}[f(X_{n+1})\mid X_{n}=x]\quad\text{and}\quad[\overline{T}f](x)=\overline{\mathbb{E}}_{\mathcal{P}_{\mathcal{T}}}[f(X_{n+1})\mid X_{n}=x],$
where $\mathcal{P}_{\mathcal{T}}$ is the set of all homogeneous Markov chains with transition matrix in $\mathcal{T}$. The following proposition establishes that these infima and suprema are always attained.
Proposition 2.16 ([19, Lemma 1]).
Let $\mathcal{T}$ be a set of transition matrices on $\mathcal{X}$. Then, for all $f\in\mathbb{R}^{\mathcal{X}}$, there exists an extreme point $T\in\mathcal{T}$ such that $Tf=\underline{T}f$, and an extreme point $T^{\prime}\in\mathcal{T}$ such that $T^{\prime}f=\overline{T}f$.
Lower and upper transition operators are conjugate, meaning that the lower transition operator $\underline{T}$ induces a corresponding upper transition operator
@@
and vice-versa. We now list some important properties of lower and upper transition operators.
Lemma 2.17 ([5, 19, 20]).
Given any lower transition operator $\underline{T}$ and its conjugate upper transition operator $\overline{T}$, for all $f,g\in\mathbb{R}^{\mathcal{X}}$ and all $n\in\mathbb{N}$ it holds that
@@
  4. 4.
  5. 5.
    $\underline{T}(\lambda f)=\lambda\underline{T}f$ and $\overline{T}(\lambda f)=\lambda\overline{T}f$ for all $\lambda\geq 0$.
The singleton set $\mathcal{T}=\{T\}$ for a certain transition matrix $T$ has $\underline{T}=\overline{T}=T$. Thus, all these properties derived for lower and upper transition operators also hold for transition matrices.
@@ 3 Characterization of Lower and Upper Hitting Probabilities
In this section, we characterise lower and upper hitting probabilities as the unique solutions of nonlinear fixed-point equations. We first formalise the notion of reachability in the imprecise setting.
@@ 3.1 Lower and Upper Reachability
Before characterising lower and upper hitting probabilities, we identify the states that can reach the target set with positive lower (resp. upper) probability. Recall that in the precise setting, for a homogeneous Markov chain with transition matrix $T$, we focused on the set of states $A^{c}\setminus\mathcal{C}_{T}$. This was done because Proposition 2.5 directly yields the hitting probabilities for other states, with the hitting probability equal to one for states in $A$ and equal to zero for states in $\mathcal{C}_{T}$. Clearly, the result for the states in $A$ carries over to the imprecise setting, as $p^{T}(x)=1$ for all $T\in\mathcal{T}$ and all $x\in A$. Hence also $\underline{p}^{\mathcal{T}}(x)=\overline{p}^{\mathcal{T}}(x)=1$ for all $x\in A$. To characterise the states with zero lower and upper hitting probability, we need suitable definitions of reachability for the imprecise setting.
Definition 3.1 (Upper reachability $\operatorname{\mathbf{(UR0)}}$).
Given a set of transition matrices $\mathcal{T}$ and a set of states $\mathcal{D}\subseteq\mathcal{X}$ we say that $x$ upper reaches $\mathcal{D}$ if there exists $n\in\mathbb{N}_{0}$ and $T\in\mathcal{T}$ such that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$. We denote by $x\rightharpoonup\mathcal{D}$ if $x$ upper reaches $\mathcal{D}$.
In the literature [5, 13, 24] another definition often appears, namely
(UR1):
$x$ upper reaches $\mathcal{D}$ if there exists $n\in\mathbb{N}_{0}$ such that $\overline{T}^{n}\mathbb{1}_{\mathcal{D}}](x)>0$.
We show that these two definitions are equivalent.
Theorem 3.2.
Definition $\operatorname{(UR0)}$ and definition $\operatorname{(UR1)}$ are equivalent.
Proof.
$\operatorname{(UR0)}\Rightarrow\operatorname{(UR1)}:$ This implication follows directly from the fact that
$[\overline{T}^{n}\mathbb{1}_{\mathcal{D}}](x)\coloneqq\sup_{T_{1},\dots,T_{n}\in\mathcal{T}}[T_{1}\cdots T_{n}\mathbb{1}_{\mathcal{D}}](x).$ (12)
Thus, if there exist $n\in\mathbb{N}_{0}$ and $\hat{T}\in\mathcal{T}$ such that $\hat{T}^{n}\mathbb{1}_{\mathcal{D}}](x)>0$ (definition (UR0)), then
$[\overline{T}^{n}\mathbb{1}_{\mathcal{D}}](x)$ $=\sup_{T_{1},\dots,T_{n}\in\mathcal{T}}[T_{1}\cdots T_{n}\mathbb{1}_{\mathcal{D}}](x)$
$\geq\sup_{T\in\mathcal{T}}[T^{n}\mathbb{1}_{\mathcal{D}}](x)$
$\geq[\hat{T}^{n}\mathbb{1}_{\mathcal{D}}](x)>0.$
$\operatorname{(UR1)}\Rightarrow\operatorname{(UR0)}:$ Now suppose that there exists $n\in\mathbb{N}_{0}$ such that $\overline{T}^{n}\mathbb{1}_{\mathcal{D}}](x)>0$. Since $\mathcal{T}$ is compact, this supremum is attained, i.e. there exist $T_{1},\dots,T_{n}\in\mathcal{T}$ such that
$T_{1}\cdots T_{n}\mathbb{1}_{\mathcal{D}}](x)=[\overline{T}^{n}\mathbb{1}_{\mathcal{D}}](x)>0.$
From this, we get
$T_{1}\cdots T_{n}\mathbb{1}_{\mathcal{D}}](x)=\sum_{\begin{subarray}{c}z_{1},\dots,z_{n-1}\in\mathcal{X}\\ y\in\mathcal{D}\end{subarray}}T_{1}(x,z_{1})\cdots T_{n}(z_{n-1},y)>0,$
hence there exists a sequence $(\hat{z}_{k})_{k=1}^{n}\coloneqq\hat{z}_{0},\hat{z}_{1},\dots,\hat{z}_{n-1},\hat{z}_{n}$, with $x=\hat{z}_{0}$ and $y=\hat{z}_{n}$, such that
$T_{1}(x,\hat{z}_{1})\cdots T_{n}(\hat{z}_{n-1},y)>0.$
We now construct a matrix $\tilde{T}$ row-wise using convexity and the SSR property. For all $z\in\mathcal{X}$ define
$\tilde{T}(z,\cdot)\coloneqq\begin{cases}\frac{1}{|\{\hat{z}_{k}=z\}|}\sum_{\hat{z}_{k}=z}T_{k+1}(\hat{z}_{k},\hat{z}_{k+1})&\text{if $z\in(\hat{z}_{k})_{k}$ }\\ S(z,\cdot)\text{ for an arbitrary $S\in\mathcal{T}$}&\text{if $z\notin(\hat{z}_{k})_{k}$}.\end{cases}$
By convexity and SSR, $\tilde{T}\in\mathcal{T}$, and $\tilde{T}(z,w)>0$ for every pair of consecutive states $z,w\in\mathcal{X}$ in the sequence $(\hat{z}_{k})_{k}$. It follows that
$[\tilde{T}^{n}\mathbb{1}_{\mathcal{D}}](x)$ $=\sum_{\begin{subarray}{c}z_{1},\dots,z_{n-1}\in\mathcal{X}\\ y\in\mathcal{D}\end{subarray}}\tilde{T}(x,z_{1})\cdots\tilde{T}(z_{n-1},y)$
$\geq\tilde{T}(x,\hat{z}_{1})\cdots\tilde{T}(\hat{z}_{n-1},y)>0,$
which is what we wanted to prove. ∎
@@
Given a set of transition matrices $\mathcal{T}$ and a set of states $\mathcal{D}\subseteq\mathcal{X}$ we say that $x$ lower reaches $\mathcal{D}$ if for all $T\in\mathcal{T}$ there exists $n\in\mathbb{N}_{0}$ such that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$. We denote by $x\rightharpoondown\mathcal{D}$ if $x$ lower reaches $\mathcal{D}$.
We present three alternative natural definitions of lower reachability that often appear in the literature [4, 24, 17, 22].
@@
and where $k^{*}$ is the first index for which $\mathcal{D}_{k^{*}}=\mathcal{D}_{k^{*}+1}$. [22, 17]
@@
$x$ lower reaches $\mathcal{D}$ if there exists $n\in\mathbb{N}_{0}$ such that for all $T\in\mathcal{T}$ it holds that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$.
@@
$x$ lower reaches $\mathcal{D}$ if $\underline{T}^{n}\mathbb{1}_{\mathcal{D}}](x)>0$ for some $n\in\mathbb{N}_{0}$. [4, 24]
One reason for this abundance of possible definitions of lower reachability is that the quantifiers do not commute, i.e. in general
@@
The set $\mathcal{E}\subseteq\mathcal{X}$ of states that lower reach $\mathcal{D}$ can be seen as
$\mathcal{E}$ $\coloneqq\{x\in\mathcal{X}\mid x\rightharpoondown\mathcal{D}\}$
$\ =\{x\in\mathcal{X}\mid\forall T\in\mathcal{T},\exists n\in\mathbb{N}_{0}:[T^{n}\mathbb{1}_{\mathcal{D}}](x)>0\}$
$\ =\bigcup_{k\in\mathbb{N}_{0}}\{x\in\mathcal{X}\mid\forall T\in\mathcal{T},\exists n\leq k:[T^{n}\mathbb{1}_{\mathcal{D}}](x)>0\}=:\bigcup_{k\in\mathbb{N}_{0}}\mathcal{E}_{k}.$
To prove the equivalence $\operatorname{(LR0)}\Leftrightarrow\operatorname{(LR1)}$ it suffices to show that $\mathcal{D}_{k}=\mathcal{E}_{k}$ for all $k\in\mathbb{N}_{0}$.
We prove this by induction on $k$. The induction base $k=0$ is trivial since $\mathcal{D}_{0}=\mathcal{D}=\mathcal{E}_{0}$. Assuming that $\mathcal{D}_{k}=\mathcal{E}_{k}$ we show that $\mathcal{D}_{k+1}=\mathcal{E}_{k+1}$, and to do so we prove the two inclusions, $\mathcal{D}_{k+1}\subseteq\mathcal{E}_{k+1}$ and $\mathcal{E}_{k+1}\subseteq\mathcal{D}_{k+1}$.
First, we observe that $(\mathcal{D}_{k})$ and $(\mathcal{E}_{k})$ are sequences of nested sets, namely $\mathcal{D}_{k}\subseteq\mathcal{D}_{k+1}$ and $\mathcal{E}_{k}\subseteq\mathcal{E}_{k+1}$. To show that $\mathcal{D}_{k+1}\subseteq\mathcal{E}_{k+1}$, we note that for all $x\in\mathcal{D}_{k+1}\setminus\mathcal{D}_{k}$, we have $\underline{T}\mathbb{1}_{\mathcal{D}_{k}}](x)>0$. Therefore, for all $T\in\mathcal{T}$ there is a state $y=y(T)\in\mathcal{D}_{k}$ such that $T(x,y)>0$, and, since $\mathcal{E}_{k}=\mathcal{D}_{k}\ni y$, there exists $n\leq k$ such that $T^{n}\mathbb{1}_{\mathcal{D}}](y)>0$. It follows that
$T^{n+1}\mathbb{1}_{\mathcal{D}}](x)\geq T(x,y)[T^{n}\mathbb{1}_{\mathcal{D}}](y)>0,$
thus $x\in\mathcal{E}_{k+1}$, proving that $\mathcal{D}_{k+1}\subseteq\mathcal{E}_{k+1}$.
To show that $\mathcal{E}_{k+1}\subseteq\mathcal{D}_{k+1}$, it suffices to prove that if $x\notin\mathcal{D}_{k+1}$ then $x\notin\mathcal{E}_{k+1}$, namely there exists $\hat{T}\in\mathcal{T}$ such that $\hat{T}^{n+1}\mathbb{1}_{\mathcal{D}}](x)=0$ for all $n\leq k$. We construct the matrix $\hat{T}$ row-wise starting from $x$: since $x\notin\mathcal{D}_{k+1}$, there exists a matrix $S_{x}\in\mathcal{T}$ such that $S_{x}\mathbb{1}_{\mathcal{D}_{k}}](x)=0$, and we set $\hat{T}(x,\cdot)=S_{x}(x,\cdot)$. Observe that
$[\hat{T}^{n+1}\mathbb{1}_{\mathcal{D}}](x)$ $=\sum_{y\in\mathcal{D}_{k}}\overset{=0}{\cancel{\hat{T}(x,y)}}[\hat{T}^{n}\mathbb{1}_{\mathcal{D}}](y)+\sum_{z\notin\mathcal{D}_{k}}\hat{T}(x,z)[\hat{T}^{n}\mathbb{1}_{\mathcal{D}}](z)$
$=\sum_{z\notin\mathcal{D}_{k}}\hat{T}(x,z)[\hat{T}^{n}\mathbb{1}_{\mathcal{D}}](z),$
@@
for all $n\leq k$. Now, for all $z\notin\mathcal{D}_{k}=\mathcal{E}_{k}$, there exists $S_{z}\in\mathcal{T}$ such that $S_{z}^{n}\mathbb{1}_{\mathcal{D}}](z)=0$ for all $n\leq k$. By setting $\hat{T}(z,\cdot)=S_{z}(z,\cdot)$, for all $z\notin\mathcal{D}_{k}$ we get
$\hat{T}^{n+1}\mathbb{1}_{\mathcal{D}}](x)=\sum_{z\notin\mathcal{D}_{k}}\hat{T}(x,z)\overset{=0}{\cancel{[\hat{T}^{n}\mathbb{1}_{\mathcal{D}}](z)}}=0,$
@@
In the rest of the paper we use (LR0) (Definition 3.3) (or the equivalent (LR1)) to define lower reachability. (LR0) is weaker than (LR2) and (LR3), as stated in the following proposition.
Proposition 3.5.
We have the following chain of implications:
$\operatorname{(LR3)}\Rightarrow\operatorname{(LR2)}\Rightarrow\operatorname{(LR1)}\Leftrightarrow\operatorname{(LR0)}.$ (13)
Proof.
The implication $\text{(LR3)}\Rightarrow\text{(LR2)}$ holds because if $\underline{T}^{n}\mathbb{1}_{\mathcal{D}}](x)>0$ for some $n\in\mathbb{N}_{0}$ then for all sequences of transition matrices $T_{1},\dots,T_{n}\in\mathcal{T}$ we have $T_{1}\cdots T_{n}\mathbb{1}_{\mathcal{D}}](x)>0$. It follows that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$ for all $T\in\mathcal{T}$.
Now assume (LR2): there exists $n^{*}\in\mathbb{N}_{0}$ such that $T^{n^{*}}\mathbb{1}_{\mathcal{D}}](x)>0$ for all $T\in\mathcal{T}$. Since (LR0) requires for all $T\in\mathcal{T}$ the existence of a $n=n(T)$ with the same property, the universal $n^{*}$ satisfies (LR0). ∎